OpenStudy (anonymous):

Using laws of exponents, simplify the following, and please show work: (zⁿ)^3/zⁿz^3

OpenStudy (lalaly):

z^3n / z^3+n = z^3n-(3+n) = z^2n-3

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

\[\frac{(z^n)^3}{z^nz^3}\]

OpenStudy (anonymous):

thats the question ↑

OpenStudy (anonymous):

= \[z^{3n - (3 + n)} = z^{2n - 3}\]